OpenStudy (anonymous):

how can i solve x^2-3x+1 using "complete the square" method?

OpenStudy (anonymous):

First, isolate the x-terms. so subtract one from both sides. Next, complete the sq. by dividing the coefficient on x (bx in the formula ax+bx+cx) by 2 and then squaring that number. whatever you come up with needs to be added to both sides. For example, y -3=x^2+4x the 4x turns to 2x when I divide it, and then that sqared is 4, so I add 4 to both sides y-3+4=x^2+2x+4 Now, simplify the equation.

OpenStudy (anonymous):

Does that make some sense? (:

OpenStudy (anonymous):

gonna read it a couple of times, lol

OpenStudy (anonymous):

okay. khanacademy.com has some great lessons on "completing the square" too @sta880

OpenStudy (anonymous):

so far i've got something like this x^2-3x+2.25=-1+2.25 , is this the right path?

OpenStudy (anonymous):

I would think so. Are you on FLVS?

OpenStudy (anonymous):

x-1.25=sqrt1.25, answer must be in significant figures, so should I just mark it as unsolveable?

OpenStudy (anonymous):

I'm not sure. @ranga can you help us?

OpenStudy (anonymous):

does this mean "solve for \(x\): \(x^2-3x+1 =0\)?

OpenStudy (anonymous):

oh, yeah, sorry

OpenStudy (anonymous):

if you want the complete the square method, first step is subtract \(1\) from both sides and write \[x^2-3x=-1\] then since half of \(3\) is \(\frac{3}{2}\) go right to \[(x-\frac{3}{2})^2=-1+(\frac{3}{2})^2=-1+\frac{9}{4}=\frac{5}{4}\]

OpenStudy (anonymous):

once you have \[(x-\frac{3}{2})^2=\frac{5}{4}\] take the square root of both sides, don't forget the \(\pm\) and write \[x-\frac{3}{2}=\frac{\sqrt5}{2}\]

OpenStudy (anonymous):

oops i forgot the \(\pm\) it should be \[x-\frac{3}{2}=\pm\frac{\sqrt5}{2}\] then finish by adding \(\frac{3}{2}\) to both sides getting \[x=\frac{3}{2}\pm\frac{\sqrt5}{2}\] or if you prefer \[x=\frac{3\pm\sqrt5}{2}\]